Question 3 Draw the AFD, SFD and BMD for the frame shown below. 10 kNm 20 KN
Answer #1

무튼 10 2OkN 10 kN/m moment about F. + EME=o, YAX10 – 10*3*(+7) – 20%2 –104550 7) v = 34,5 KN forces Summation of all verticalSummation of all horizontal forces at Fx=0, H = 10 kN. Analyze the members of the Isolating the members and frame separately​​​​​☆ To LOKN Member AB Axial Force - locarlon Ankar we have to balance T3m the other han end. I 34.5 kn 405 kn essing IFy 20 undBending moment AT A, BMA=0 At B, BM= 34.5 % -103* = 58.5 km .. 58. Som z BMD for AB. Member Be 1.4.5kn it lour, 58.5 unm 15msf at B = -10 kn Sf at c = +10+10=okr. A 3 (SFO)BC loten BM at B = 58.5 urm BM at ca 58.5-10x5 28.5 krm 8.5km 21 (BMD) Best XSF at c = 4.5 kn St Just left of D = 4.5 UN sf just night of D=4.5-20 = -15.5 un Sf at E= -15.5415.5=0 kara 405 KN € (SAD) CABM at c = 8.5 kerm BM at D = 8.5 44.545=gl korm BM at E= 8.5 44.547 -20x2 = 0 krim sl urm go5 + (Bano DecMember Ef 15.5 不Ek1buy 15.5kn Axial Force = 15.5 kr (compressives 1515 kn (AFD) EF 18.5kn sf at Ero Sf at fao I (SED) EF BM a